

********************************************************************
UARS WEEKLY STATUS REPORT - 16 AUG 2002
********************************************************************
***********
SOLSTICE
***********
Nothing received.
***********
SUSIM
***********
SUSIM successfully performed its daily mid-resolution and MgII
solar scans during periods of predicted UARS telemetry.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
PEM continues to work with the GUVI team on the diagnosis of data from
the first conjunction period. Analysis of the two conjunction passes showed
that the aurora as detected by GUVI was at higher latitude than UARS.
GUVI images from other times from this conjunction period show spatial
regions that both GUVI and PEM sampled when aurora was present. Separation
is several minutes to 10's of minutes. A PEM representative attended
the TIMED workshop August 7-9.
PEM data from the second GUVI-PEM conjunction is now being processed.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered. NOAA has changed the binary structure of the NMC data and
PEM is now receiving two data streams, one in the original VAX format and the
other from their new binary format.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations to support re-engineering of the UARS
Mission Operations Concept to develop capabilities for the
Science Traceability Mission.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 03 August 2002 (Orbit 59572, GMT Day 215) through
09 August 2002 (Orbit 59677, GMT Day 221).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
SOLSTICE 217/1325 59610 POWER ON
SUSIM 217/1649 59612 POWER ON
PEM 218/1423 59626 AXIS-1 POWER ON
PEM 218/1504 59627 AXIS-1 POWER OFF/ON
HALOE 220/1320 59655 POWER ON
HALOE 220/1320 59655 POWER OFF
HALOE 220/1501 59657 POWER ON
PEM 221/1318 59670 POWER ON - ZEPS/HEPS
DAY ONLY
The beta angle ranged between the angles of 1.31 degrees to 27.2
degrees for this report period. The beta angle is now increasing
towards a maximum angle of 45 degrees on 19 August 2002 (DOY 231).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 1.29 to 1.21 degrees C. Additional
battery performance data for each battery on 02 August and 09 August
2002 are provided in Appendix D.
The clock error ranged from -10.2 to 10.8 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 07
SOLSTICE = 02
SUSIM = 02
PEM = 02
HALOE = 02
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 01
Two (2) Anomaly Reports remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
One (1) operational support problem occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The Raytheon development team has delivered a new database patch which
should fix the status buffer display problem. This patch has been
tested this week on the prototype system but the FOT is waiting for
some more fixes before it will be installed on the operational system.
The FOT continues to work with the WINDII instrument team to determine
the source of their instrument turn on problem. It was recently
discovered, by the WINDII team, that WINDII's science data has not
been captured correctly since February. There were discrepancies
found in their POCC resident microloads during Eclipse transition and
the investigation has been focused on determining if these loads are
correct. An engineering test will be done on 10 August to try to
isolate the source of the problem.
MEETINGS
The FOT attended the UARS MSR on 06 August 2002.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 04 September 2002 (DOY 247).
This maneuver will be a Reverse to Forward Flight maneuver (Southern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 09 August 2002
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
59580 215/1244 02-268 MPS YAW (return to VT Level 5)
59581 215/1426 02-269 MPS BATSOC RESET
59610 217/1241 02-270 SOLSTICE MUB ENABLE
59610 217/1325 02-271 SOLSTICE POWER ON
59612 217/1649 02-272 SUSIM POWER ON (PART-1)
59613 217/1757 02-272 SUSIM POWER ON (PART-1 COMPLETE)
59623 218/1000 02-273 OBC/CLK CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
59626 218/1423 02-274 PEM AXIS-1 POWER ON
59627 218/1504 02-274 PEM AXIS-1 POWER OFF/ON
59640 219/1238 02-272 SUSIM POWER ON COMPLETE
59640 219/1238 02-276 SUSIM SUSIM CMDS
59641 219/1423 02-275 OBC/CLK CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
59643 219/1754 02-277 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
59655 220/1320 02-278 HALOE POWER ON
59655 220/1320 02-279 HALOE POWER OFF
59657 220/1501 02-278 HALOE POWER ON
59658 220/1647 02-280 OBC/CLK CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
59664 221/0301 02-282 OBC/CLK CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
59670 221/1318 02-283 PEM POWER ON - ZEPS/HEPS DAY-ONLY
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 09 August 2002
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
______________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 09 August 2002
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
59623 218/1000 CLK ADJ FR 0.00 TO -20.6 MSEC/DAY
59641 219/1433 CLK ADJ FR -20.6 TO 0.00 MSEC/DAY
59643 219/1839 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
59658 220/1656 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
59664 221/0305 CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 09 August 2002
02 August 2002 - GMT Day 214 Beta = 0.0 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -157.0/-168.0 -3.05/-4.62 +00.0/+00.0 08.5 0.000 0.0
2 +11.2/-69.6 +5.47/+4.08 +27.0/-13.4 23.8 1.090 20.4
3 +22.4/-50.4 +4.51/+2.92 +25.7/-13.5 23.8 1.080 20.0
09 August 2002 - GMT Day 221 Beta = 27.2 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -146.0/-168.0 -3.05/-4.01 +00.0/+00.0 08.5 0.000 0.0
2 +11.2/-123.0 +5.79/+4.53 +27.2/-17.4 27.5 1.030 21.0
3 +22.4/-44.8 +5.08/+3.87 +26.4/-17.4 27.5 1.020 20.5
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
59580 215/1248 L6 -> L5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 09 August 2002
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
02-044 FOT 59629 218/1756 31261 Equipment Fail-over at WS
50 second data loss, non-recov
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
02-041 FOT 59279 195/1156 N/A TMON 01 tripped
02-042 FOT 59503 210/0959 N/A TMON 01 tripped
02-043 FOT 59571 214/2308 31245 Data Drop Out
02-044 FOT 59629 218/1756 31261 Equipment Fail-over at WS
50 second data loss, non-recov
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
02-039 FOT See List July 30832 Generic Late Acquisition
02-044 FOT 59629 218/1756 31261 Equipment Fail-over at WS
50 second data loss, non-recov